F(x)= √(Tanx) + √(sinx)
1
Expert's answer
2020-06-01T19:00:50-0400
"f(x)=\sqrt{\tan(x)}+\sqrt{\sin(x)}"

"{df \over dx}={1 \over 2\sqrt{\tan(x)}}({1 \over \cos^2x})+{1 \over 2\sqrt{\sin(x)}}(\cos(x))"